What is the dilution factor of the examples below:

           

            a) 20 uL serum added to 180 uL of saline

            b) 50 uL of serum added to 450 uL of saline

            c) 50 uL of serum added to 200 uL of saline

            d) 100 uL of serum added to 900 uL of saline

            e) 25 uL of serum added to 475 uL of saline

            f) 50 uL of serum added to 150 uL of saline

            g) 5 uL of serum added to 195 uL of saline
Answer #1

Formula for dilution factor, DF is:

DF = (Final volume) / (Solute volume)

(a): Volume of solute = 20 uL

Final volume = 20 uL + 180 uL = 200 uL

DF = (Final volume) / (Solute volume) = 200 uL / 20 uL = 10 (Answer)

(b): Volume of solute = 50 uL

Final volume = 50 uL + 450 uL = 500 uL

DF = (Final volume) / (Solute volume) = 500 uL / 50 uL = 10 (Answer)

(c): Volume of solute = 50 uL

Final volume = 50 uL + 200 uL = 250 uL

DF = (Final volume) / (Solute volume) = 250 uL / 50 uL = 5 (Answer)

(d): Volume of solute = 100 uL

Final volume = 100 uL + 900 uL = 1000 uL

DF = (Final volume) / (Solute volume) = 1000 uL / 100 uL = 10 (Answer)

(e): Volume of solute = 25 uL

Final volume = 25 uL + 475 uL = 500 uL

DF = (Final volume) / (Solute volume) = 500 uL / 25 uL = 20 (Answer)

(f): Volume of solute = 50 uL

Final volume = 50 uL + 150 uL = 200 uL

DF = (Final volume) / (Solute volume) = 200 uL / 50 uL = 4 (Answer)

(g): Volume of solute = 5 uL

Final volume = 5 uL + 195 uL = 200 uL

DF = (Final volume) / (Solute volume) = 200 uL / 5 uL = 40 (Answer)